Find two level curve of the function: f: R^2→R defined by : f(x,y)= 4x^2+ 16y^2+8

Draw their rough sketches


1
Expert's answer
2021-04-14T13:11:03-0400
"f(x,y)=4x^2+16y^2+8"


"z=4x^2+16y^2+8"

"z=c"


"c=8:"


"4x^2+16y^2+8=8"


"4x^2+16y^2=0"


"Point (0,0)"

"c=16:"


"4x^2+16y^2+8=16"


"4x^2+16y^2=8"


"\dfrac{x^2}{(\sqrt{2})^2}+\dfrac{y^2}{(\dfrac{\sqrt{2}}{2})^2}=1, ellipse"


"c=24:"


"4x^2+16y^2+8=24"


"4x^2+16y^2=16"


"\dfrac{x^2}{(2)^2}+\dfrac{y^2}{(1)^2}=1, ellipse"